OpenStudy (anonymous):

Gavin's boat took 2 hours to make a trip downstream with a 4-mph current. The return trip against the same current took 3 hours. Find the speed of the boat in still water.

OpenStudy (bahrom7893):

satellite can u go back and check my work?

OpenStudy (anonymous):

got this one too?

OpenStudy (bahrom7893):

well yea lol

OpenStudy (bahrom7893):

if u want me to haha

OpenStudy (anonymous):

go nuts

OpenStudy (bahrom7893):

uhmm right let's organize this: let d be the distance traveled.

OpenStudy (bahrom7893):

Let V be the speed of the boat, then: V+4 = d/2 V-4 = d/3

OpenStudy (bahrom7893):

now solve for V

OpenStudy (anonymous):

okay for the first one I got v=-4+2d? I am anywhere close?

OpenStudy (bahrom7893):

no it's d/2 not 2d. V = -4+d/2. Now plug that into the 2nd equation

OpenStudy (anonymous):

okay I'm stuck I got d/2-8=d/3 I don't know where to go from there